# Multi-Metric Preference Alignment for Generative Speech Restoration

### v1

@@ 4.4 Application: Empowering Discriminative Models via Pseudo-Labeling
To showcase the practical utility of our aligned models, we explore their potential as “data annotators” in a scenario where ground-truth is scarce: real-world singing voice restoration. We took a standard discriminative restoration model, Voicefixer, and fine-tuned it on a set of singing recordings. The key difference is that the “clean” supervision signals were not ground-truth recordings, but pseudo-labels generated by our DPO-enhanced AnyEnhance model.
@@
As shown in Table 6, the fine-tuned Voicefixer model achieves dramatic improvements across all metrics compared to its original state. This result is significant: it demonstrates that a powerful, DPO-aligned generative model can effectively create high-quality supervision signals, enabling smaller, more efficient discriminative models to be trained for tasks where obtaining paired data is prohibitively difficult. This bridges the gap between the two modeling paradigms and opens up new avenues for practical speech restoration. More detail on this application can be found in the Appendix D.4.
@@ A.1 DPO for Autoregressive Models
The derivation for AR models follows the original DPO framework (Rafailov et al. 2023). The goal is to optimize a policy $p_{\theta}(y|x)$ to maximize a reward function $r(x,y)$, while regularizing its deviation from a reference policy $p_{\text{ref}}(y|x)$ using a KL-divergence term. The RL objective is:
@@ A.3 DPO for Flow-Matching Models
Because FM training assumes a Gaussian likelihood, following practice in (Wallace et al. 2024; Liu et al. 2025), we can express the policies into velocity predictions $v_{\theta}$ and $v_{\text{ref}}$ as follows:
@@ B.1 Model Details
Masked Generative Model (MGM). We use AnyEnhance as our model for MGM paradigm. AnyEnhance (Zhang et al. 2025a) is a unified masked generative model for speech and singing voice restoration, capable of handling a wide range of tasks (e.g., denoising, dereverberation, super-resolution) via prompt-guided in-context learning and a self-critic sampling mechanism. It is based on a transformer architecture (Vaswani et al. 2017) trained with a masked language modeling objective, allowing the model to predict clean audio tokens of DAC codec (Kumar et al. 2024) conditioned on the degraded mel-spectrogram. We use the pre-trained checkpoint in the original work, which comprises $\sim$360M parameters and is trained on a large-scale dataset comprising 20k hours of speech and singing data.
@@
Autoregressive (AR) Model. Our model for AR paradigm, which we term AR+Soundstorm, follows a two-stage pipeline inspired by the architecture introduced in MaskGCT (Wang et al. 2025b; Zhang et al. 2024c). The first stage employs a 0.5B parameter autoregressive model, initialized from the weights of Qwen2.5-0.5B-Instruct [^3] https://huggingface.co/Qwen/Qwen2.5-0.5B-Instruct. This model takes the noisy audio’s dense features from a pre-trained w2v-bert-2 model [^4] https://huggingface.co/facebook/w2v-bert-2.0 at the 17th-layer as input and is trained to predict the semantic tokens of clean audio extracted from MaskGCT’s semantic codec [^5] https://huggingface.co/amphion/MaskGCT/tree/main/semantic_codec. We use MaskGCT’s second stage [^6] https://huggingface.co/amphion/MaskGCT/tree/main/s2a_model, which is a non-autoregressive Soundstorm-based model (Borsos et al. 2023) with approximately 300M parameters. This model converts the predicted semantic tokens into acoustic tokens, which are then synthesized into the final clean audio waveform.
@@
Flow-Matching (FM) Model. Our FM model, named Flow-SR, is designed to learn a continuous mapping from degraded to clean audio representations. It follows the Diffusion Transformer (DiT) (Peebles and Xie 2023) architecture, which adapts transformer-based models for continuous-time generative tasks. The core of the model consists of 10 LLaMA-2-style decoder layers with 1024 hidden dimensions and 16 attention heads, comprising approximately 220M parameters. The model is also conditioned on dense features of noisy audios from the 17th layer of the pre-trained w2v-bert-2 model and is trained using a flow-matching objective (Lipman et al. 2022) to predict the clean mel-spectrogram (at a 24kHz sampling rate). For the final waveform synthesis, we employ a pre-trained Vocos vocoder (Siuzdak 2023), specifically leveraging the open-source Vevo1.5 implementation [^7] https://huggingface.co/amphion/Vevo1.5/tree/main/acoustic_modeling/Vocoder (Zhang et al. 2025c; Li et al. 2025c).
B.2 Training Details
@@
Pre-Training Data: For AR+Soundstorm and Flow-SR, we sampled 7k hours of clean speech from the Emilia dataset (He et al. 2024). And we use the same noise & rir introduced in AnyEnhance (Zhang et al. 2025a) which contains $\sim$1.2k hours noise and 62,668 rir files, to generate the noisy and reverberant audios.
@@
Pre-Training Procedure: We both train AR+Soundstorm and Flow-SR using the Adam (Kingma and Ba 2014) optimizer, with a learning rate of 5e-4 and 1e-4 respectively and 32K warmup steps, following the inverse square root learning schedule. The training is performed on 4 GPUs, with each GPU processing 5000 frames of audio per step, which is equivalent to 100 seconds of audio, for 200k steps. The training takes approximately 2.5 days for both two models.
@@ C.1 Baseline Models
  • Voicefixer (Liu et al. 2022): A discriminative general speech restoration model, consists of a ResUNet-based analysis stage which predicts a ratio mask over the degraded mel-spectrogram, and a synthesis stage which converts the masked mel-spectrogram into the final waveform using TFGAN-based vocoder. We use its pre-trained checkpoint [^8] https://github.com/haoheliu/voicefixer.
  • MaskSR (Li, Wang, and Liu 2024): A masked generative model for speech restoration tasks, which predicts acoustic tokens from DAC codec (Kumar et al. 2024) conditioned on the noisy mel-spectrogram. We report the results in the original paper.
  • GenSE (Yao et al. 2025a): A two-stage autoregressive generative framework for speech enhancement, which decouples denoising and generation by first mapping noisy speech to clean semantic tokens and then generating clean acoustic tokens. We report the results in the original paper.
  • LLaSE-G1 (Kang et al. 2025): A LLaMA-based generative speech enhancement model that preserves acoustic consistency by predicting fine-grained speech tokens from continuous WavLM features, and supports diverse SE tasks like denoise, echo cancellation, etc. We report the results in the original paper.
  • FlowSE (Wang et al. 2025d): A flow-matching-based speech enhancement model that learns a continuous mapping from noisy mel-spectrogram to clean mel-spectrogram, with optional text input conditioning capability. We report the results in the original paper.
@@ C.2 Evaluation Details
  • DNSMOS (Reddy, Gopal, and Cutler 2022): a reference-free metric for 16 kHz audio, providing SIG, BAK, and OVRL scores (1-5) for signal quality, background noise, and overall quality.
  • NISQA (Mittag et al. 2021): a reference-free metric for 48 kHz audio that outputs a single overall perceptual quality score (1-5).
  • SpeechBERTScore (Saeki et al. 2024): evaluates semantic similarity between enhanced and reference signals using HuBERT-base [^9] https://huggingface.co/facebook/hubert-base-ls960.
@@
  • General Speech Restoration (GSR): To evaluate performance on complex, simultaneous distortions (e.g., noise, reverberation, clipping, and bandwidth limitation), we use three test sets: the official Voicefixer-GSR dataset (Liu et al. 2022), and two simulated sets introduced in AnyEnhance based on clean speech from LibriVox (Kearns 2014) and clean singing from CCMusic (Zhou et al. 2024).
  • Speech Enhancement (SE): To evaluate performance specifically on downstream SE tasks, we use the official test sets from the DNS Challenge 2020 (Reddy et al. 2020), including both the with-reverb and no-reverb conditions.
@@ D.4 Application: Empowering Discriminative Models via Pseudo-Labeling
Pseudo-Label Generation and Fine-tuning. We first prepared a specialized AnyEnhance model, which was pre-trained with additional singing voice data via the SingNet pipeline (Gu et al. 2025) and then aligned on singing voice data with our multi-metric preference alignment strategy. This expert generative model was then used to process the noisy recordings from our training set, generating high-quality pseudo-clean labels for each segment.

### v2

@@ 4.4 Application: Empowering Discriminative Models via Pseudo-Labeling
We demonstrate our aligned model’s utility as a “data annotator” in a data-scarce scenario by fine-tuning the Voicefixer restoration model on singing recordings. Instead of ground-truth data, we used pseudo-labels generated by our DPO-enhanced AnyEnhance model as supervision. This resulted in dramatic performance improvements across all metrics (Table 6). Our findings show a powerful generative model can effectively create supervision signals to train smaller, discriminative models when paired data is prohibitively difficult to obtain. This bridges the gap between the two modeling paradigms and opens up new avenues for practical speech restoration. More detail on this application can be found in the Appendix D.4.
@@ 5 Conclusion
Acknowledgments
This work is funded by the NSFC under Grant 62376237, Shenzhen Science and Technology Program ZDSYS20230626091302006, 2023 Shenzhen stability Science Program, Internal Project Fund from Shenzhen Research Institute of Big Data, Grant No. T00120230002, and Program for Guangdong Introducing Innovative and Enterpreneurial Teams, Grant No. 2023ZT10X044.
@@ A.1 DPO for Autoregressive Models
The derivation for AR models follows the original DPO framework (rafailov2023direct). The goal is to optimize a policy $p_{\theta}(y|x)$ to maximize a reward function $r(x,y)$, while regularizing its deviation from a reference policy $p_{\text{ref}}(y|x)$ using a KL-divergence term. The RL objective is:
@@ A.3 DPO for Flow-Matching Models
Because FM training assumes a Gaussian likelihood, following practice in (wallace2024diffusiondpo; liu2025videoalign), we can express the policies into velocity predictions $v_{\theta}$ and $v_{\text{ref}}$ as follows:
@@ B.1 Model Details
Masked Generative Model (MGM). We use AnyEnhance as our model for MGM paradigm. AnyEnhance (zhang2025anyenhance) is a unified masked generative model for speech and singing voice restoration, capable of handling a wide range of tasks (e.g., denoising, dereverberation, super-resolution) via prompt-guided in-context learning and a self-critic sampling mechanism. It is based on a transformer architecture (vaswani2017attention) trained with a masked language modeling objective, allowing the model to predict clean audio tokens of DAC codec (kumar2024high) conditioned on the degraded mel-spectrogram. We use the pre-trained checkpoint in the original work, which comprises $\sim$360M parameters and is trained on a large-scale dataset comprising 20k hours of speech and singing data.
@@
Autoregressive (AR) Model. Our model for AR paradigm, which we term AR+Soundstorm, follows a two-stage pipeline inspired by the architecture introduced in MaskGCT (wang2024maskgct; amphion). The first stage employs a 0.5B parameter autoregressive model, initialized from the weights of Qwen2.5-0.5B-Instruct [^2] https://huggingface.co/Qwen/Qwen2.5-0.5B-Instruct. This model takes the noisy audio’s dense features from a pre-trained w2v-bert-2 model [^3] https://huggingface.co/facebook/w2v-bert-2.0 at the 17th-layer as input and is trained to predict the semantic tokens of clean audio extracted from MaskGCT’s semantic codec [^4] https://huggingface.co/amphion/MaskGCT/tree/main/semantic_codec. We use MaskGCT’s second stage [^5] https://huggingface.co/amphion/MaskGCT/tree/main/s2a_model, which is a non-autoregressive Soundstorm-based model (borsos2023soundstorm) with approximately 300M parameters. This model converts the predicted semantic tokens into acoustic tokens, which are then synthesized into the final clean audio waveform.
@@
Flow-Matching (FM) Model. Our FM model, named Flow-SR, is designed to learn a continuous mapping from degraded to clean audio representations. It follows the Diffusion Transformer (DiT) (peebles2023scalable) architecture, which adapts transformer-based models for continuous-time generative tasks. The core of the model consists of 10 LLaMA-2-style decoder layers with 1024 hidden dimensions and 16 attention heads, comprising approximately 220M parameters. The model is also conditioned on dense features of noisy audios from the 17th layer of the pre-trained w2v-bert-2 model and is trained using a flow-matching objective (lipman2022flow) to predict the clean mel-spectrogram (at a 24kHz sampling rate). For the final waveform synthesis, we employ a pre-trained Vocos vocoder (siuzdak2023vocos), specifically leveraging the open-source Vevo1.5 implementation [^6] https://huggingface.co/amphion/Vevo1.5/tree/main/acoustic_modeling/Vocoder (zhang2025vevo; amphion_v0.2).
B.2 Preference Dataset Construction
To construct our GenSR-Pref dataset, we first generated a diverse pool of candidate outputs from the three base models. This diversity is crucial for creating meaningful preference pairs and ensuring the robustness of the subsequent DPO training, we detail the specific stochastic sampling configurations used for each model paradigm below:
  • MGM: We varied the number of decoding steps by randomly sampling from the range [10, 40]. We also employed nucleus sampling with a fractional top-k value of 0.9 and a temperature of 1.0 to encourage diversity.
  • AR: For the autoregressive model, we used a combination of sampling techniques: top-k sampling with k=25, nucleus sampling (top-p) with p=0.8, and a temperature of 1.0.
  • FM: We introduced stochasticity by varying the number of sampling steps (randomly chosen from [10, 40]) and by using different Classifier-Free Guidance (CFG) scales, which were randomly sampled from the range [1.0, 3.0].
B.3 Training Details
@@
Pre-Training Data: For AR+Soundstorm and Flow-SR, we sampled 7k hours of clean speech from the Emilia dataset (he2024emilia). And we use the same noise & rir introduced in AnyEnhance (zhang2025anyenhance) which contains $\sim$1.2k hours noise and 62,668 rir files, to generate the noisy and reverberant audios.
@@
Pre-Training Procedure: We both train AR+Soundstorm and Flow-SR using the Adam (kingma2014adam) optimizer, with a learning rate of 5e-4 and 1e-4 respectively and 32K warmup steps, following the inverse square root learning schedule. The training is performed on 4 GPUs, with each GPU processing 5000 frames of audio per step, which is equivalent to 100 seconds of audio, for 200k steps. The training takes approximately 2.5 days for both two models.
@@ C.1 Baseline Models
  • Voicefixer (liu2022voicefixer): A discriminative general speech restoration model, consists of a ResUNet-based analysis stage which predicts a ratio mask over the degraded mel-spectrogram, and a synthesis stage which converts the masked mel-spectrogram into the final waveform using TFGAN-based vocoder. We use its pre-trained checkpoint [^7] https://github.com/haoheliu/voicefixer.
  • MaskSR (li2024masksr): A masked generative model for speech restoration tasks, which predicts acoustic tokens from DAC codec (kumar2024high) conditioned on the noisy mel-spectrogram. We report the results in the original paper.
  • GenSE (anonymous2025gense): A two-stage autoregressive generative framework for speech enhancement, which decouples denoising and generation by first mapping noisy speech to clean semantic tokens and then generating clean acoustic tokens. We report the results in the original paper.
  • LLaSE-G1 (kang2025llase): A LLaMA-based generative speech enhancement model that preserves acoustic consistency by predicting fine-grained speech tokens from continuous WavLM features, and supports diverse SE tasks like denoise, echo cancellation, etc. We report the results in the original paper.
  • FlowSE (wang2025flowse): A flow-matching-based speech enhancement model that learns a continuous mapping from noisy mel-spectrogram to clean mel-spectrogram, with optional text input conditioning capability. We report the results in the original paper.
@@ C.2 Evaluation Details
  • DNSMOS (reddy2022dnsmos): a reference-free metric for 16 kHz audio, providing SIG, BAK, and OVRL scores (1-5) for signal quality, background noise, and overall quality.
  • NISQA (mittag2021nisqa): a reference-free metric for 48 kHz audio that outputs a single overall perceptual quality score (1-5).
  • SpeechBERTScore (saeki2024speechbertscore): evaluates semantic similarity between enhanced and reference signals using HuBERT-base [^8] https://huggingface.co/facebook/hubert-base-ls960.
@@
  • General Speech Restoration (GSR): To evaluate performance on complex, simultaneous distortions (e.g., noise, reverberation, clipping, and bandwidth limitation), we use three test sets: the official Voicefixer-GSR dataset (liu2022voicefixer), and two simulated sets introduced in AnyEnhance based on clean speech from LibriVox (kearns2014librivox) and clean singing from CCMusic (ccmusic).
  • Speech Enhancement (SE): To evaluate performance specifically on downstream SE tasks, we use the official test sets from the DNS Challenge 2020 (reddy2020interspeech), including both the with-reverb and no-reverb conditions.
@@ D.4 Application: Empowering Discriminative Models via Pseudo-Labeling
Pseudo-Label Generation and Fine-tuning. We first prepared a specialized AnyEnhance model, which was pre-trained with additional singing voice data via the SingNet pipeline (gu2025singnet) and then aligned on singing voice data with our multi-metric preference alignment strategy. This expert generative model was then used to process the noisy recordings from our training set, generating high-quality pseudo-clean labels for each segment.